Roman want to determine the volume of three oddly shaped bottles. all he has is gallon container filled with water. After much experimenting, he discovers the following facts. the gallon container fills the first and second containers. the first and the third containers exactly fill the second container. three times the volume of the first container fills the second. what is the volume of each container?

Thank You!!

Answer by macston(5194)

F=first container; S=second container; T=third container
1)F+S=1 gal
2)F+T=S
3)3F=S
F+3F=1 gal=4F 1F=1/4 gal ANSWER 1: The first container holds 1 quart.
S=3F=3/4 gal ANSWER 2: The second container holds 3 quarts.
F+T=S
1/4gal+T=3/4gal
T=1/2 gal ANSWER 3: The third container holds 2 quarts.